# Hyperbolic Fuzzy $C$-Means with Adaptive Weight-based Filtering for Clustering in Non-Euclidean Spaces

### v1

@@ 3.3 P_SFCM
$\mathbf{x}_{i},\mathbf{v}_{j}\in\mathbb{D}_{\alpha}^{p},\forall i\in[n],j\in[c].$ (19)
6.1 Membership Update
The membership weights $w_{ij}$ are updated as:
$w_{ij}=\frac{d_{H}(\mathbf{x}_{i},\mathbf{v}_{j})^{-2/(m-1)}}{\sum_{l=1}^{c}d_{H}(\mathbf{x}_{i},\mathbf{v}_{l})^{-2/(m-1)}}.$ (20)
6.2 Centroid Update
The hyperbolic centroid update is computed using logarithmic and exponential maps on the Poincaré ball:
$\mathbf{v}_{j}^{(t+1)}=\exp_{\mathbf{v}_{j}^{(t)}}\left(\frac{\sum_{i=1}^{n}(w_{ij})^{m}\cdot\log_{\mathbf{v}_{j}^{(t)}}(\mathbf{x}_{i})}{\sum_{i=1}^{n}(w_{ij})^{m}}\right),$ (21)
where, $\log_{\mathbf{v}}(\cdot)$ and $\exp_{\mathbf{v}}(\cdot)$ are the logarithmic and exponential maps as described in 8, 9.
@@ 6.2 Centroid Update
(Bounded distance) For any $\mathbf{x},\mathbf{y}\in\mathbb{D}_{\alpha}^{p}$: $0\leq d_{H}(\mathbf{x},\mathbf{y})<\infty$.
$d_{H}\geq 0$ follows from the properties of $\tanh^{-1}$ and the norm in Equation 7. $d_{H}(x,y)$ is finite, since $\|-\mathbf{x}\oplus_{\alpha}\mathbf{y}\|<\frac{1}{\sqrt{\alpha}}$ in $\mathbb{D}_{\alpha}^{p}$. Detailed proof of this lemma is discussed in the supplementary.
@@
$f(\mathbf{v}_{j})=\sum_{i=1}^{n}(w_{ij})^{m}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}).$
Proof of this lemma is given in the supplementary document.
@@
We have shown the convergence analysis of the proposed objective function 19.
Step 1: Descent Property of Membership Update
We aim to show that the HypeFCM algorithm converges to a stationary point of the following objective function defined on the Poincaré ball model $\mathbb{D}_{\alpha}^{p}$: For fixed $\mathbf{V}^{(t)}$, the weight update minimizes $J_{m}$ subject to constraints:
$w_{ij}^{(t+1)}=\frac{d_{H}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})^{-2/(m-1)}}{\sum_{k=1}^{c}d_{H}(\mathbf{x}_{i},\mathbf{v}_{k}^{(t)})^{-2/(m-1)}}.$
For each point $i$, form the Lagrangian from Equation 19.
$L_{i}(w_{i1},...,w_{ic},\lambda_{i})=\sum_{j=1}^{c}(w_{ij})^{m}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})+\lambda_{i}(\sum_{j=1}^{c}w_{ij}-1).$ (22)
@@
$\frac{\partial{L_{i}}}{\partial{w_{ij}}}=m(w_{ij})^{m-1}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})+\lambda_{i}=0.$
@@
$m(w_{ij})^{m-1}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})+\lambda_{i}=0$
$\implies w_{ij}=(-\frac{\lambda_{i}}{md_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})})^{1/(m-1)}.$
@@
$\sum_{j=1}^{c}(-\frac{\lambda_{i}}{md_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})})^{1/(m-1)}=1.$
@@
$\alpha_{i}\sum_{j=1}^{c}(d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)}))^{-1/(m-1)}=1.$
@@
$\alpha_{i}=\frac{1}{\sum_{j=1}^{c}(d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)}))^{-1/(m-1)}}.$
@@
$w_{ij}^{(t+1)}=\frac{d_{H}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})^{-2/(m-1)}}{\sum_{j=1}^{c}d_{H}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})^{-2/(m-1)}}.$
@@
Step 2: Descent Property of Centroid Update
@@
$\mathbf{v}_{j}^{(t+1)}=\exp_{\mathbf{v}_{j}^{(t)}}\left(\frac{\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}\log_{\mathbf{v}_{j}^{(t)}}(\mathbf{x}_{i})}{\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}}\right),$ (23)
where $\log_{\mathbf{v}}(\cdot)$ and $\exp_{\mathbf{v}}(\cdot)$ are the logarithmic and exponential maps at $\mathbf{v}$ in the Poincaré Disc. model. This update minimizes the cost function with respect to $\mathbf{v}_{j}$ under Riemannian geometry, which conforms with the one in Equation 17. For fixed $\mathbf{W}^{(t+1)}$, the centroid update minimizes
$J_{m}:\mathbf{v}_{j}^{(t+1)}=\operatorname*{argmin}_{\mathbf{v}_{j}\in\mathbb{D}_{\alpha}^{p}}\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}).$
@@
Step 3: Monotonicity and Convergence
$\{J_{m}^{(t)}\}$ is monotonically decreasing sequence in t by 24 and $J_{m}^{(t)}\geq 0$ by Lemma 1. Therefore, $\{J_{m}^{(t)}\}$ converges to some $J^{*}\geq 0$, by Monotone Convergence theorem [38].
@@
Therefore, the unfiltered objective acts as a pointwise majorizer to the one with filtration.
Under these conditions, and by properties of alternating minimization on manifolds [39], the sequence of updates converges to a stationary point of the objective function.
@@
7.1 Details of the Datasets.
We validate the efficacy of HypeFCM on some synthetic and real-world datasets. Iris, Glass, Ecoli, Wine, Phishing URL, Abalone, Zoo, ORHD (Optical Recognition of Handwritten Digits) datasets are taken from UCI machine learning repository [40]; Flights, MNIST datasets are taken from Kaggle, and cure-t1-2000n-2D, cure-t2-4k, donutcurves, disk-4000n, smile1, 3MC are taken from Clustering Benchmark datasets available at https://github.com/deric/clustering-benchmark. Experiments on some of these datasets are shown in the supplementary.
7.2 Experimental Setup & Baselines.
We carried out experiments on HypeFCM on $6$ synthetic datasets along with $7$ real-world datasets. The performance of HypeFCM is measured by involving two well-known performance metrics, namely Adjusted Rand Index (ARI) [41] and Normalized Mutual Information (NMI) [42].
The notable benchmark methods, namely the PCM [13], FCM [15], P_SFCM [26], $k$-means [1], MinibatchKmeans [43], and HSFC [27] clustering methods are considered for comparison with our proposed HypeFCM. A brief description of the datasets used for experimentation is available in Table III.
Figure 2: Grid search plots of $2$ real-world datasets (a) Ecoli, (b) Wine and $2$ synthetic datasets (c) Smile1, (d) Cure-t1-2000n, for two ranges of curvatures ($\alpha$), 00-$1$ and $1$-$1000$ and different filtration ($k$) values against the performance metric, NMI are presented above.
7.3 Parameter Setting.
@@
We have performed two separate $3$D grid search experiments for each dataset with two parameters (curvature ($\alpha$), filtration value ($k$)) against the performance metric, NMI with one varying the curvature ($\alpha$) from 00 to $1$ by increasing the step size of $0.1$ and the other from $1$ to $1000$ by increasing the step size of $100$ in each iteration on two real-world datasets and two synthetic datasets. The grid search plots in Figure 2 show the performance of HypeFCM on two real-world datasets; Ecoli, Wine, and two synthetic datasets; smile1, cure-t1-2000n-2D respectively.
We also performed two separate $2$D plots for NMI with varying curvature ($\alpha$), for the values of filtration ($k$) varying from $1$ to $15$. We plotted them in two separate plots for $k=1,5,10,15$ in figure 3 & 4 for two real datasets: Ecoli, Wine; and two synthetic datasets: smile1, cure-t1-2000n-2D respectively. We analyze the performance from the figures and conclude that not much variation in performance is observed with the choice of curvature ($\alpha$) regarding the best choice of filtration value ($k$).
@@
We have applied HypeFCM on $6$ synthetic datasets. Performances are presented in Table I. Our HypeFCM algorithm underscores almost all contenders in terms of ARI and NMI. The Average Rank in Table I is just the average of the ranks of ARI, and NMI values obtained from the executions of each method for $15$ times, all initialized with the same seeding at a time. We calculate the $p$-values from the Wilcoxon signed-rank test [44] for the methods compared to HypeFCM and represent them in a table in the supplementary document. Smaller $p$-values ($\leq 0.05$) indicate a statistically significant difference between the performances of the corresponding algorithm and HypeFCM, denoted by or else if there is no statistically significant difference, denoted by .
7.5 Experiments on Real Benchmark Datasets.
We applied HypeFCM on $7$ real-world benchmark datasets and the performances are presented in Table II. Beyond the superior performance in synthetic datasets, we also achieved notable performances on real-world datasets as well. We have demonstrated the t-SNE visualization [45] of HypeFCM in Figure 1 for three real-world datasets, Glass, Ecoli, Wine respectively. The Average Rank mentioned in the table is the same as above for all the real-world datasets. denotes a statistically significant difference between the performances of the corresponding algorithm with our proposed HypeFCM, and indicates no significant difference.
Figure 4: The performance metric NMI vs curvature ($\alpha$) plots of HypeFCM (ours) for two synthetic datasets (a) Smile1, (b) Cure-t1-2000n-2D respectively.

### v2

@@ 6 Convergence Analysis of HypeFCM
$\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}\in\mathbb{D}_{\alpha}^{p},\forall i\in[n],j\in[c].$ (18)
Membership Update. The membership weights $w_{ij}$ are updated as:
$w_{ij}=\frac{d_{hyp}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j})^{-2/(m-1)}}{\sum_{l=1}^{c}d_{hyp}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{l})^{-2/(m-1)}}.$ (19)
Centroid Update. The hyperbolic centroid update is computed using logarithmic and exponential maps on the Poincaré ball.
$\mathbf{v}_{j}^{(t+1)}=\exp_{\mathbf{v}_{j}^{(t)}}\left(\frac{\sum_{i=1}^{n}(w_{ij})^{m}\cdot\log_{\mathbf{v}_{j}^{(t)}}(\mathbf{x}_{i}^{\prime})}{\sum_{i=1}^{n}(w_{ij})^{m}}\right),$ (20)
where, $\log_{\mathbf{v}}(\cdot)$ and $\exp_{\mathbf{v}}(\cdot)$ are the logarithmic and exponential maps as described in 5, 6. We show mathematically that this centroid updation belongs to the Poincaré ball. The centroids $\{\mathbf{v}^{(t)}_{j}\}_{j=1}^{c}$ using the Riemannian log-exp map as Equation 16. Let,
$\delta:=\tanh\left(\frac{\sqrt{\alpha}\lambda_{\mathbf{v}_{j}}\|\bar{v}_{j}\|}{2}\right).$
The hyperbolic tangent satisfies the following:
$\tanh(z)\in(0,1),\quad\text{for }z>0.$
In our case,
$z=\frac{\sqrt{\alpha}\lambda_{\mathbf{v}_{j}}\|\bar{v}_{j}\|}{2}>0,$
since: $\alpha>0$ implies $\sqrt{\alpha}>0$, $\lambda_{\mathbf{v}_{j}}=\frac{2}{1-\alpha\|\mathbf{v}_{j}\|^{2}}>0$ as $\|\mathbf{v}_{j}\|<\frac{1}{\sqrt{\alpha}}$, $\|\bar{v}_{j}\|\geq 0$, hence,
$\delta=\tanh\left(\frac{\sqrt{\alpha}\lambda_{\mathbf{v}_{j}}\|\bar{v}_{j}\|}{2}\right)<1.$
Thus, $\delta<1\implies\|\frac{\delta}{\sqrt{\alpha}}\cdot\frac{\bar{v}_{j}}{\|\bar{v}_{j}\|}\|<\frac{1}{\sqrt{\alpha}}$. Thus by Lemma 1, $\mathbf{v}_{j}^{(t)}\in\mathbb{D}_{\alpha}^{p}$. This shows that the centroid update belongs to the Poincaré disc.
@@
For any $\mathbf{x},\mathbf{y}\in\mathbb{D}_{\alpha}^{p}$: $0\leq d_{hyp}(\mathbf{x},\mathbf{y})<\infty$.
The detailed proof of this lemma is discussed in the Appendix A.
Figure 2: Grid search plots of $2$ real-world datasets (a) Ecoli, (b) Wine and $2$ synthetic datasets (c) Smile1, (d) Cure-t1-2000n, for two ranges of curvatures ($\alpha$), 0-11 and 11-$1000$ and different filtration ($k$) values against the performance metric, NMI are presented above.
@@
$f(\mathbf{v}_{j})=\sum_{i=1}^{n}(w_{ij})^{m}d_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}).$
The proof of this lemma is discussed in Appendix A.
@@
We have proved step by step the convergence analysis of the proposed objective function 18.
Step 1: Descent Property of Membership Update.
We aim to show that the HypeFCM algorithm converges to a stationary point of the following objective function defined on the Poincaré ball model $\mathbb{D}_{\alpha}^{p}$: For fixed $\mathbf{V}^{(t)}$, the weight update minimizes $J_{m}$ subject to constraints in Equation 18.
For each point $i$, form the Lagrangian
$L_{i}(w_{i1},...,w_{ic},\lambda_{i})=\sum_{j=1}^{c}(w_{ij})^{m}d_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}^{(t)})+\lambda_{i}(\sum_{j=1}^{c}w_{ij}-1).$ (21)
@@
$\frac{\partial{L_{i}}}{\partial{w_{ij}}}=m(w_{ij})^{m-1}d_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}^{(t)})+\lambda_{i}=0.$
@@
$m(w_{ij})^{m-1}d_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}^{(t)})+\lambda_{i}=0$
$\implies w_{ij}=(-\frac{\lambda_{i}}{md_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}^{(t)})})^{1/(m-1)}.$
@@
$\sum_{j=1}^{c}(-\frac{\lambda_{i}}{md_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}^{(t)})})^{1/(m-1)}=1.$
@@
$\alpha_{i}\sum_{j=1}^{c}(d_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}^{(t)}))^{-1/(m-1)}=1.$
@@
$\alpha_{i}=\frac{1}{\sum_{j=1}^{c}(d_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}^{(t)}))^{-1/(m-1)}}.$
@@
$w_{ij}^{(t+1)}=\frac{d_{hyp}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}^{(t)})^{-2/(m-1)}}{\sum_{j=1}^{c}d_{hyp}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}^{(t)})^{-2/(m-1)}}.$
@@
Step 2: Descent Property of Centroid Update.
@@
$\mathbf{v}_{j}^{(t+1)}=\exp_{\mathbf{v}_{j}^{(t)}}\left(\frac{\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}\log_{\mathbf{v}_{j}^{(t)}}(\mathbf{x}_{i}^{\prime})}{\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}}\right),$ (22)
where $\log_{\mathbf{v}}(\cdot)$ and $\exp_{\mathbf{v}}(\cdot)$ are the logarithmic and exponential maps at $\mathbf{v}$ in the Poincaré Disc model. This update minimizes the cost function concerning $\mathbf{v}_{j}$ under Riemannian geometry, which conforms with the one in Equation 16. For fixed $\mathbf{W}^{(t+1)}$, the centroid update minimizes
$J_{m}:\mathbf{v}_{j}^{(t+1)}=\operatorname*{argmin}_{\mathbf{v}_{j}\in\mathbb{D}_{\alpha}^{p}}\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}d_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}).$
@@
Step 3: Monotonicity and Convergence.
$\{J_{m}^{(t)}\}$ is monotonically decreasing sequence in t by Equation 23 and $J_{m}^{(t)}\geq 0$ by Lemma 1. Therefore, $\{J_{m}^{(t)}\}$ converges to some $J^{*}\geq 0$, by Monotone Convergence theorem [42].
@@
Therefore, the unfiltered objective acts as a pointwise majorizer to the one with filtration. Under these conditions, and by properties of alternating minimization on manifolds [43], the sequence of updates converges to a stationary point of the objective function. ∎
@@
7.1 Details of the Datasets
We validate the efficacy of HypeFCM on $6$ synthetic and $12$ real-world datasets. Iris, Glass, Ecoli, Wine, Wisconsin B.C., Phishing URL, Abalone, Glass, Zoo, ORHD (Optical Recognition of Handwritten Digits) datasets are taken from UCI machine learning repository [44]; Flights, MNIST datasets are taken from the Kaggle, and Cure-t1-2000n-2D, Cure-t2-4k, Donutcurves, Disk-4000n, Smile1, 3MC are taken from Clustering Benchmark datasets available at
https://github.com/deric/clustering-benchmark. Experiments on some of these datasets are shown in Appendix B.
7.2 Experimental Setup & Baselines
The performance of HypeFCM is measured by involving two well-known performance metrics, namely Adjusted Rand Index (ARI) [45] and Normalized Mutual Information (NMI) [46]. The notable base methods, namely the PCM [13], FCM [15], $k$-means [1], MinibatchKmeans [47], along with the State-of-the-Art methods like P_SFCM [26], EFKM [41], UFCER [48], FCSR [40], IFKMHC [49], HSFC [27], FCPFM [29] clustering methods are considered for comparison with our proposed HypeFCM.
7.3 Parameter Setting
@@
We have performed two separate 33D grid search experiments for each dataset with two parameters (curvature ($\alpha$), filtration value ($k$) against the performance metric, NMI with one varying the curvature ($\alpha$) from 0 to $1$ by increasing the step size of $0.1$ and the other from $1$ to $1000$ by increasing the step size of $100$ in each iteration on two real-world and two synthetic datasets. The grid search plots in Figure 2 show the performance of HypeFCM on two real-world datasets, Ecoli, Wine, and two synthetic datasets, Smile1, Cure-t1-2000n, respectively.
We also performed two separate 22D plots for NMI with varying curvature ($\alpha$), for the values of filtration ($k$) varying from $1$ to $15$. We plotted them in two separate plots for $k=1,5,10,15$ in figure 3 for two real datasets: Ecoli, Wine, respectively. We analyze the performance from the figures and conclude that not much variation in performance is observed with the choice of curvature ($\alpha$) regarding the best choice of filtration value ($k$).
@@ 7.4 Experiment on Datasets
We carried out experiments on $6$ synthetic datasets and $12$ real-world datasets. Performances on some of them are presented in Table II and the rest of them in the appendix. Our HypeFCM algorithm underscores almost all contenders in terms of ARI and NMI. We have demonstrated the t-SNE visualization [50] of HypeFCM in Figure 1 for two real-world datasets, Glass, and Wisconsin Breast Cancer. The Average Rank in Table II is just the average of the ranks of ARI and NMI values obtained from the executions of each method for $15$ times, all initialized with the same seeding at a time. We calculate the $p$-values from the Wilcoxon signed-rank test [51] for the methods compared to HypeFCM and represent them in a table in the appendix. Smaller $p$-values ($\leq 0.05$) indicate a statistically significant difference between the performances of the corresponding algorithm and HypeFCM, denoted by or else denotes there is no statistically significant difference.
7.5 Ablation Studies
We conducted an ablation study by varying the key parameter, Filtration.
HypeFCM with and without Filtration. We conducted two separate experiments, one with filtration and the other without using any filtration on two real-world datasets, Wine, Ecoli, and two synthetic datasets, Cure-t1-2000n and Smile1, respectively. We have performed these experiments by setting the curvature parameter $(\alpha)$ as $1$. Figure 4 shows that using filtration improves the performance of HypeFCM compared to the same without filtration.
Figure 4: Comparison of Clustering Performance, NMI vs $k$ with and without filtration for two real-world (a) Wine, (b) Ecoli, and two synthetic (c) Cure-t1-2000n-2D, (d) Smile1 datasets respectively.
8 Discussions.
The performance results of HypeFCM demonstrate its effectiveness across both synthetic and real-world datasets. The algorithm’s superior performance in terms of ARI and NMI metrics against competing methods validates the fundamental hypothesis that hyperbolic geometry can better capture complex hierarchical relationships in clustering tasks. The significant performance of HypeFCM on $12$ real-world benchmark datasets, as visualized through t-SNE for Glass, Wisconsin B.C. datasets, demonstrates how effectively HypeFCM preserves cluster boundaries and captures the inherent hierarchical structures that often characterize real-world datasets. Integrating hyperbolic geometry via the Poincaré Disc model represents a significant methodological advancement. The model’s exponentially expanding space toward its boundary naturally accommodates hierarchical structures that are prevalent in many real-world datasets but challenging to represent in Euclidean space. This geometric advantage, combined with fuzzy clustering principles, allows HypeFCM to capture subtle hierarchical relationships that traditional fuzzy clustering methods cannot properly identify.
9 Conclusion & Future works.
Our proposed HypeFCM marks a significant advancement in handling complex hierarchical data structures by seamlessly integrating hyperbolic geometry with the principles of fuzzy clustering. Our method exploits the unique properties of the Poincaré Disc model, which provides exponentially more space towards its boundary, making it naturally suited for representing hierarchical relationships. Incorporating a weight-based filtering mechanism enhances the algorithm’s ability to focus on relevant geometric relationships while pruning less significant connections, thereby improving both computational efficiency and clustering accuracy. The algorithm’s iterative optimization process, combining hyperbolic distance calculations with fuzzy membership updates, demonstrates robust performance in capturing complex structures that traditional Euclidean-based methods struggle to represent adequately. The experimental results validate the effectiveness of our approach, showing superior performance in preserving hierarchical relationships while maintaining computational feasibility. While the current results are promising, expanding this framework to handle streaming data and exploring integration with deep learning architectures may further enhance the algorithm’s applicability across diverse domains where hierarchical relationships are intrinsic to the data structure.
Appendix A Proofs of Lemmas and Theorems from Section 6.
Proof of the Lemma 1.
Proof.
The Möbius addition of $\mathbf{x}$ and $\mathbf{y}$ in $\mathbb{D}_{\alpha}^{p}$ is defined as :
$\mathbf{x}\oplus_{\alpha}\mathbf{y}:=\frac{\left(1+2\alpha\langle\mathbf{x},\mathbf{y}\rangle+\alpha\|\mathbf{y}\|^{2}\right)\mathbf{x}+\left(1-\alpha\|\mathbf{x}\|^{2}\right)\mathbf{y}}{1+2\alpha\langle\mathbf{x},\mathbf{y}\rangle+\alpha^{2}\|\mathbf{x}\|^{2}\|\mathbf{y}\|^{2}}.$ (24)
$d_{hyp}(\mathbf{x},\mathbf{y})=(2/\sqrt{\alpha})\tanh^{-1}\left(\sqrt{\alpha}\left\|-\mathbf{x}\oplus_{\alpha}\mathbf{y}\right\|\right).$ (25)
$d_{hyp}\geq 0$ follows from the properties of $\tanh^{-1}$ and the norm in Equation 25. $d_{hyp}(\mathbf{x},\mathbf{y})$ is finite, since $\|\mathbf{x}\oplus_{\alpha}\mathbf{y}\|<\frac{1}{\sqrt{\alpha}},\forall\mathbf{x},\mathbf{y}\in\mathbb{D}_{\alpha}^{p}$ & $\|\mathbf{x}\|=\|\mathbf{-x}\|,\forall\mathbf{x}\in\mathbb{D}_{\alpha}^{p}.$
By the Cauchy-Schwarz inequality,
$\;|\langle\mathbf{x},\mathbf{y}\rangle|\leq\|\mathbf{x}\|\|\mathbf{y}\|$
$\implies$ $|\alpha\langle\mathbf{x},\mathbf{y}\rangle|\leq\alpha\|\mathbf{x}\|\|\mathbf{y}\|$
$\implies$ $\big{|}(1+2\alpha\langle\mathbf{x},\mathbf{y}\rangle+\alpha\|\mathbf{y}\|^{2})\mathbf{x}\big{|}$
$\quad<\frac{1}{\sqrt{\alpha}}\big{|}(1+2\alpha\|\mathbf{x}\|\|\mathbf{y}\|+\alpha\|\mathbf{y}\|^{2})\big{|}$
$\implies$ $\big{|}(1+2\alpha\langle\mathbf{x},\mathbf{y}\rangle+\alpha\|\mathbf{y}\|^{2})\mathbf{x}+(1-\alpha\|\mathbf{x}\|^{2})\mathbf{y}\big{|}$
$\quad<\big{|}\frac{1}{\sqrt{\alpha}}(1+2\alpha\|\mathbf{x}\|\|\mathbf{y}\|+\alpha\|\mathbf{y}\|^{2})\big{|}$
$\implies$ $\big{|}(1+2\alpha\langle\mathbf{x},\mathbf{y}\rangle+\alpha\|\mathbf{y}\|^{2})\mathbf{x}$
$\quad+(1-\alpha\|\mathbf{x}\|^{2})\mathbf{y}\big{|}<3\frac{1}{\sqrt{\alpha}}$
$\implies$ $\bigg{|}\frac{(1+2\alpha\langle\mathbf{x},\mathbf{y}\rangle+\alpha\|\mathbf{y}\|^{2})\mathbf{x}}{1+2\alpha\langle\mathbf{x},\mathbf{y}\rangle+\alpha^{2}\|\mathbf{x}\|^{2}\|\mathbf{y}\|^{2}}$
$\quad+\frac{(1-\alpha\|\mathbf{x}\|^{2})\mathbf{y}}{1+2\alpha\langle\mathbf{x},\mathbf{y}\rangle+\alpha^{2}\|\mathbf{x}\|^{2}\|\mathbf{y}\|^{2}}\bigg{|}$
$\quad<\frac{3(1/\sqrt{\alpha})}{3}$
$\implies$ $\|\mathbf{x}\oplus_{\alpha}\mathbf{y}\|<\frac{1}{\sqrt{\alpha}}.$
Figure 5: Image segmentation of natural images by different methods (a) Original Image, (b) PCM, (c) FCM, (d) EFKM, (e) HSFC, (f) HypeFCM (ours).
Figure 6: The t-SNE visualization of $5$ real-world benchmark datasets: Flights (5k), Wine, ORHD, Zoo, Phishing URL (5k) for FCM and HypeFCM clustering methods,respectively [from top to bottom].
Lemma 3.
The Poincaré disc model $\mathbb{D}_{\alpha}^{p}$ with the distance metric $d_{hyp}$ is complete.
Proof.
This follows from the Hopf-Rinow theorem [52] as the hyperbolic space is a geodesically complete Riemannian manifold. ∎